OpenStudy (anonymous):

Find the number of positive integers factors (excluding 1 and the number itself) of each of the following numbers: a)3960 b)98010

OpenStudy (anonymous):

For a) I got 6 and for b) i got 8. Can someone check if those are right?

OpenStudy (anonymous):

for a) 2*2*2*3*3*5*11=3960 b) 2*3*3*3*3*5*11*11=98010

ganeshie8 (ganeshie8):

how did you get 6 for a ?

OpenStudy (anonymous):

well there are three 2's, two 3's, one 5, and one 11, total of 6 factors for b) there is one 2, four 3's, one 5, and two 11's, total of 8 factors?

ganeshie8 (ganeshie8):

a factor is a number that divides the number evenly

OpenStudy (anonymous):

oh my bad 7 factrs

ganeshie8 (ganeshie8):

`n = a*b` `a` and `b` are factors of `n`

ganeshie8 (ganeshie8):

2*2*2*3*3*5*11=3960 Notice that 2 is a factor of 3960 and so is 2*2

ganeshie8 (ganeshie8):

2*2*2*3 is also a factor

ganeshie8 (ganeshie8):

you need to consider all the combinations

OpenStudy (anonymous):

so yea, for a) there are 7 factors and for b) there are 8

ganeshie8 (ganeshie8):

can you list all the 7 factors for a ?

ganeshie8 (ganeshie8):

you should get more than 7 for a

OpenStudy (anonymous):

2x1980 2x2x990 2x2x2x495 2x2x2x3x165 2x2x2x3x3x55 2x2x2x3x3x5x11 <--7 factors

OpenStudy (ikram002p):

u just need to know number of factors theorem -.-

ganeshie8 (ganeshie8):

i dont see any factors in your list however all those are same as 3960

OpenStudy (ikram002p):

is it in ur txt book ?

OpenStudy (anonymous):

This is handout for my teacher, not in the textbook

ganeshie8 (ganeshie8):

here is a list of all factors including 1 and the number itself http://www.wolframalpha.com/input/?i=divisors+3960

OpenStudy (anonymous):

so wait would these be the factors 2, 3, 4, 5, 6, 8, 9, 10, 11, 12, 15, 18, 20, 22, 24, 30, 33, 36, 40, 44, 45, 55, 60, 66, 72, 88, 90, 99, 110, 120, 132, 165, 180, 198, 220, 264, 330, 360, 396, 440, 495, 660, 792, 990, 1320, 1980, ooooh okay, i get it, i have to list all the multiples

OpenStudy (ikram002p):

-.-

ganeshie8 (ganeshie8):

you don't need to list all the factors, you just need to give them the total number of factors

OpenStudy (anonymous):

oh okay, i get it, thanks guys

OpenStudy (anonymous):

So for the first there are 11 factors right?

ganeshie8 (ganeshie8):

how 11 ?

OpenStudy (anonymous):

oh my bad 48 factors

ganeshie8 (ganeshie8):

that includes 1 and 3960

OpenStudy (anonymous):

so 46 then

ganeshie8 (ganeshie8):

but they don't want them in the list so subtract 2

ganeshie8 (ganeshie8):

yes!

ganeshie8 (ganeshie8):

but how can you find the count without using wolfram ?

OpenStudy (anonymous):

that i don't know

OpenStudy (anonymous):

can you teach me?

ganeshie8 (ganeshie8):

its easy

ganeshie8 (ganeshie8):

whats the prime factorization of 3960 ?

OpenStudy (anonymous):

2*2*2*3*3*5*11

ganeshie8 (ganeshie8):

write it in exponent form

OpenStudy (anonymous):

(2^3)(3^2)(5^1)(11^1)

ganeshie8 (ganeshie8):

\[\large 3960 = 2^{\color{Red}{3}}3^{\color{Red}{2}}5^{\color{Red}{1}}11^{\color{red}{1}}\]

OpenStudy (anonymous):

right

ganeshie8 (ganeshie8):

now consider a factor, it can have an exponent of 2 in 0,1,2,3 ways right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

it can have an exponent of 3 in 0,1,2 ways

ganeshie8 (ganeshie8):

it can have an exponent of 5 in 0,1 ways

ganeshie8 (ganeshie8):

it can have an exponent of 11 in 0,1 ways

OpenStudy (anonymous):

oh okay, so when you multiply the choices you get 48

ganeshie8 (ganeshie8):

so the number of ways of choosing the exponents of prime numbers in a divisor is : \[\large \color{Red}{(3+1)(2+1)(1+1)(1+1)}\]

ganeshie8 (ganeshie8):

exactly!

OpenStudy (anonymous):

okay i get it, thats what i was missing

ganeshie8 (ganeshie8):

try to find the number of factors of number in part b using combinations

OpenStudy (anonymous):

98010=(2^1)(3^4)(5^1)(11^2)  there are 60 factors of 98010 right?

ganeshie8 (ganeshie8):

Looks perfect! not that you need to subtract 2 factors since the questions asks you to exclude 1 and the number itself

ganeshie8 (ganeshie8):

the count 60 includes both 1 and 98010 : \(\large 1 = 2^03^05^011^0\) \(\large 98010 = 2^13^45^111^3\)

OpenStudy (anonymous):

oh okay so you subtract 2

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

Thanks for the help everyone :D